OpenStudy (anonymous):

a rectangular plot of land has a perimeter of 40m. if the area of the plot is at least 75 m, what is the narrowest that this plot of land can be?

OpenStudy (anonymous):

what is the widest?

OpenStudy (anonymous):

i got (x-15)(x-5)>=0 is this correct? what is the answer?

OpenStudy (anonymous):

@eta

OpenStudy (anonymous):

See: you are given perimeter : \[2l + 2w = 40 \\ l + w = 20\]

OpenStudy (anonymous):

i got xy>=75 and 2x+2y=40 as my 2 equations

OpenStudy (anonymous):

i was taught to solve for 2x+2y=40 and then plug it into the inequality

OpenStudy (anonymous):

yes these two equations are right..

OpenStudy (anonymous):

then i got (x-15)(x-5)>=0

OpenStudy (anonymous):

yes, do the same..

OpenStudy (anonymous):

but what is the answer?

OpenStudy (anonymous):

i found that its (inf,5] and [15,inf)

OpenStudy (anonymous):

How you got that?

OpenStudy (anonymous):

I mean : \(x + y = 20\), right?

OpenStudy (anonymous):

So, \(y = 20 - x\)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):